Two rays with common end points `O` forms a 30 degree angle. Point A lies on one ray, point B on the other ray and AB=1. The maximum possible length of OB is and options are (A) = root 3 (B) = 4/root 3 (C) = (root 3 +1)/root 2 (D) = 2

Grade:12

1 Answers

Arun
6 years ago
Dear Meet
 
Triangle OAB has the property that angle O = 30° and AB = 1
From sine law
(sin A)/OB = (sin O)/AB
(sin A)/OB = (1/2)/1
OB = 2 sinA
Mazimum of sinA is 1 when A = 90°>
Hence maximum of OB = 2
Option D is correct.
Regards
Arun (askIITians forum expert)